OpenStudy (anonymous):

PLEASE HELP?? reward an medal will be granted 9.) what multiplies to = -40 and adds/subtracts to equal -3??? 10.) What multiplies to equal -60 and adds/subtracts to equal 7???? 11.) what multiplies to equal 24 and adds/subtracts to equal -10????

OpenStudy (anonymous):

@mathmale @nincompoop @Shamil98 @Wolfe8 @charlotte123 @Nuke_ur_friend @Hero @adrynicoleb PLEASE HELP

OpenStudy (wolfe8):

9) xy=-40, x(-/+)y=-3 10) xy=-60, x(-/+)y=7 11) xy=24, x(-/+)y=-10 I think that's how you set it up

OpenStudy (anonymous):

9.) x^2 -3x-40=0 10.) x^2 +7x-60=0 11.) x^2-10x+24=0 12.) x^2-12x+20=0 14.) x^2-17x+70=0 X-box factoring last numeral goes on top part of x and middle x number goes on bottom. I'm struggling findng what numbers multiply together to equal top and add together to equal bottom though????? @wolfe8

OpenStudy (anonymous):

@Elliot943

OpenStudy (anonymous):

@Mel98

OpenStudy (anonymous):

@bibby

OpenStudy (anonymous):

I don't know how to do this, but I tagged Bibby, he may be able to

OpenStudy (anonymous):

Question is this where we draw an 'x' and multiple the \(x^2\) term by the constant and then put it on top of the 'x'? x^2 -3x-40=0 So: -3 on the bottom of 'x'. 1 * -40 = -40 on the top of 'x'.

OpenStudy (anonymous):

@LanguageEnthusiast I have that part done/ But what two numbers multiply to equal -40 and add/subtract to equal -3??

OpenStudy (anonymous):

-8 * 5 = -40 -8 + 5 = -3

OpenStudy (anonymous):

10) -5 * 12 = -60 -5 + 12 = 7

OpenStudy (anonymous):

@perl please look above i asked like 5 questions

OpenStudy (anonymous):

11) -6 * -4 = 24 -6 + (-4) = 10

OpenStudy (anonymous):

12) -10 * -2 = 20 -10 + (-2) = -12

OpenStudy (perl):

you can make a table of factors,

OpenStudy (anonymous):

yeah i did that i just need help with the 2 numbers. tho that's what stumps me.

OpenStudy (perl):

pick one you want to work on,

OpenStudy (perl):

and make it a tough one ;)

OpenStudy (perl):

just kidding :D

OpenStudy (anonymous):

14.) x^2-17x+70=0 15.) x^2-13x+12=0 16.) x^2+15x+44=0 Is 3 too many????

OpenStudy (perl):

i can only solve one at a time , ok lets do 14)

OpenStudy (perl):

brb, one sec , afk

OpenStudy (perl):